Simplify: 3^2 × 10^4 Simplify: 3^2 × 10^4 Simplify: 3^2 × 10^4

Simplify:  3^2 × 10^4
Simplify:  3^2 × 10^4
Simplify:  3^2 × 10^4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The above question can be written as,
= 3 × 3 × 10 × 10 × 10 × 10 = 9 × 10000
= 90000

Thanks